# MI9: An Integrated Runtime Governance Framework for Agentic AI

### v3

@@ 2 Related Work
Alignment research has primarily focused on training-time interventions: Reinforcement Learning from Human Feedback (RLHF) [christiano_deep_2017, ouyang_training_2022] and Constitutional AI [anthropic_constitutional_ai_2022] establish initial value alignment but cannot address failures emerging during autonomous operation when agents encounter novel situations or coordinate with other systems [kenton_alignment_2021].
Building on this foundation, recent work has mapped agentic system taxonomies [schneider_generative_2025, kasirzadeh_characterizing_2025], governance frameworks [openai_whitepaper_2024, raza_trism_2025, engin_dimensional_2025, kolt_governing_2025], and threat models [narajala_securing_2025, chan_atfaa_2025, syros_saga_2025]. However, leading benchmarks prioritize task completion over governance dimensions such as behavioral consistency [kapoor_agents_that_matter_2024, liu_survey_agent_evaluation_2025, zhou_webarena_2023, jimenez_swe_bench_2023, sumers_agentverse_2025].
Meanwhile, current monitoring solutions [wu_agentops_2024, langfuse_observability_2024, langsmith_tracing_2024, wandb_llm_monitoring_2024, datadog_llm_2024] provide reactive observation rather than proactive intervention. Similarly, process observability research [fournier_process_2025] and visibility frameworks [chan_visibility_2024] focus on observation, while enterprise platforms [holisticai_governance_2024, monitaur_platform_2024, modelop_agentic_2024] and security frameworks [owasp_llm_top10_2024, nist_ai_rmf_2024] rely on static risk assessment inadequate for emergent runtime behaviors.
@@ 4.1 Framework Integration and Overview
This integrated architecture enables proactive, real-time oversight specifically designed for agentic systems exhibiting emergent behaviors during execution, addressing the fundamental gap between static pre-deployment assessments and reactive post-incident analysis. Production deployments require standard distributed systems coordination [fowler_event_2005, bailis_coordination_2014], but the core governance semantics operate independently of the underlying consistency mechanisms.
@@ 4.3 Agentic Telemetry Schema
Traditional infrastructure monitoring captures operational events (HTTP responses, database latency, service calls) but systematically misses the cognitive processes that create governance risks in agentic systems. The majority of agentic governance violations originate from cognitive behaviors such as goal revision, memory retrieval, tool-chaining decisions—that remain invisible to conventional observability frameworks [fournier_process_2025]. Safe deployment of agentic AI systems requires visibility into the moments when agents autonomously revise objectives, chain unexpected tool sequences, or retrieve memory that fundamentally alters downstream behavior—cognitive processes critical for responsible oversight yet absent from standard infrastructure telemetry.
@@
ATS extends OpenTelemetry’s emerging agent conventions [opentelemetry_genai_2024] by introducing governance-semantic abstractions that transform opaque agent execution into actionable oversight intelligence. Policy engines evaluate event attributes to enforce constraints, such as "Tier 2 agents cannot execute shell commands without approval," while drift detectors analyze cognitive event patterns to identify concerning behavioral changes. This semantic foundation enables the real-time intervention capabilities that reactive monitoring lacks: in governance terms, we cannot govern what we cannot observe.
@@ 4.5 Real-Time Conformance Engine
Building on formal runtime verification frameworks for adaptive systems [carwehl_runtime_2023], we implement a sequence-aware rule layer operating on the ATS stream. Our approach employs finite-state machines (FSMs) following proven Communicating Sequential Processes (CSP) verification principles [luckcuck_varanus_2024], where each rule compiles into an FSM with states representing pattern progress and transitions triggered by ATS events that satisfy specified predicates. This design balances expressiveness with computational efficiency: FSMs encode sequential and temporal constraints relevant to agent governance while maintaining bounded memory usage and deterministic evaluation with $O(k)$ event processing time per agent, where $k$ is the number of active patterns.
@@ 4.6 Behavioral Drift Detection
Building on recent advances in uncertainty-based drift detection [komorniczak_drift_2024] and multi-agent ensemble methods [vieira_driftage_2021], our system analyzes behavioral patterns using statistical evaluation of ATS event sequences within defined goal contexts. It maintains frequency distributions of event types, tool usage, and sequence characteristics for each distinct agent objective, constructing goal-conditioned baselines $B_{g}$ that enable detection when behavior changes while goals remain constant.
@@ 7 Acknowledgments
8 Appendix
8.1 Agency-Risk Index Scoring Criteria
The following tables present the comprehensive scoring criteria used to evaluate AI agent capabilities across three fundamental dimensions: autonomy, adaptability, and continuity. Each dimension is assessed using multiple criteria with a standardized 0–3 scoring scale.
Table 6: Autonomy Dimension Scoring Criteria
Criterion Scoring Scale (0–3)
Multi-step Planning 0: Single-action responses only
1: Sequential 2–3 step plans with fixed logic
2: Multi-step plans (4+ steps) with conditional branching or parallel execution
3: Recursive planning with dynamic replanning and sub-goal generation
Goal Management 0: Fixed goals, no modification capability
1: Parameter adjustment within predefined goal boundaries
2: Goal refinement and prioritization based on environmental feedback
3: Autonomous goal creation, modification, and objective redefinition
Resource Orchestration 0: Single tool/resource per task
1: Sequential tool usage following predefined workflows
2: Parallel resource coordination with dependency management
3: Dynamic resource discovery, delegation to subagents, novel workflow creation
Decision Authority 0: Human approval required for all actions
1: Autonomous for routine tasks, approval for resource allocation or external actions
2: Self-directed within defined parameters, escalates only exceptional cases
3: Full decision autonomy with post-hoc reporting and accountability
Table 7: Adaptability Dimension Scoring Criteria
Criterion Scoring Scale (0–3)
Strategy Evolution 0: Fixed strategy, no modification capability
1: Parameter tuning within existing strategy framework
2: Switching between predefined strategies based on performance metrics
3: Novel strategy synthesis through experimentation and analysis
Performance Learning 0: No learning from outcomes, static responses
1: Simple pattern recognition and response adjustment
2: Systematic improvement from success/failure feedback within task domain
3: Meta-learning across domains with knowledge transfer and generalization
Environmental Responsiveness 0: Operates only in single, predefined context
1: Predefined responses to known environmental variations
2: Dynamic behavioral adjustment to changing conditions
3: Adaptive responses to novel environments with context inference
Interaction Adaptation 0: Fixed interaction patterns regardless of counterpart
1: Limited variation in communication style within role constraints
2: Behavioral modulation based on counterpart type and task requirements
3: Dynamic personality and communication style adaptation
Table 8: Continuity Dimension Scoring Criteria
Criterion Scoring Scale (0–3)
Memory Architecture 0: No memory retention between interactions
1: Session-based memory (retains context within single session)
2: Persistent memory with selective retention and updates
3: Hierarchical memory with forgetting mechanisms and knowledge consolidation
Operational Continuity 0: Restarts fresh each interaction, no context carryover
1: Basic context preservation between related interactions
2: Multi-session continuity with relationship and preference tracking
3: Long-term operational persistence across extended timeframes
State Complexity 0: Stateless operation, no internal state tracking
1: Basic state variables for current task progress
2: Multiple concurrent context management with state synchronization
3: Hierarchical state management with predictive state preparation
Knowledge Integration 0: No knowledge accumulation across interactions
1: Retains frequently used patterns and standard procedures
2: Cross-task knowledge transfer and experience accumulation
3: Meta-cognitive knowledge integration with conceptual abstraction
8.2 Agency-Risk Index (ARI) Calculation
The Agency-Risk Index provides a quantitative assessment of an AI agent’s governance requirements based on its inherent capabilities across three fundamental dimensions of agency. The ARI serves as the foundational risk classification that drives containment threshold determination and governance intensity scaling.
8.2.1 Mathematical Formulation
The ARI aggregates capability assessments across autonomy, adaptability, and continuity dimensions using equal weighting to reflect the distinct governance challenges posed by each dimension:
$\text{ARI}=\frac{1}{3}\sum_{d=1}^{3}\left(\frac{1}{12}\sum_{c=1}^{4}s_{d,c}\right)$ (1)
where:
  • $d\in\{1,2,3\}$ represents the three agency dimensions (Autonomy, Adaptability, Continuity)
  • $c\in\{1,2,3,4\}$ represents the four criteria within each dimension
  • $s_{d,c}\in\{0,1,2,3\}$ is the capability score for criterion $c$ in dimension $d$
  • $\text{ARI}\in[0,1]$ provides a normalized risk assessment
8.2.2 Dimensional Scoring and Aggregation
Each dimension contributes equally to the final ARI calculation, reflecting the principle that autonomy, adaptability, and continuity represent orthogonal governance challenges requiring distinct oversight mechanisms.
Dimension-Level Calculation:
For each dimension $d$, the dimensional score is computed as:
$D_{d}=\frac{1}{4}\sum_{c=1}^{4}\frac{s_{d,c}}{3}$ (2)
Final ARI Computation:
$\text{ARI}=\frac{1}{3}(D_{\text{autonomy}}+D_{\text{adaptability}}+D_{\text{continuity}})$ (3)
8.2.3 Risk Tier Classification
[…]
8.3 Risk-Tier-Weighted Containment Model Development
MI9’s containment framework requires organizations to develop context-sensitive escalation models that balance operational continuity with risk mitigation. This section provides detailed guidance for constructing organization-specific containment policies.
8.3.1 Comprehensive Containment Decision Framework
$\text{Containment Level}=\operatorname{argmax}_{c\in C}P(c|\text{Risk Tier},\text{Context},\text{Policy})$ (5)
where $C=\{\text{Monitor},\text{Planning},\text{Restriction},\text{Isolation}\}$.
8.3.2 Multi-Dimensional Context Assessment
Table 9: Context Factor Scoring Framework Example
Context Dimension Low Risk (1) Medium Risk (2) High Risk (3)
Business Impact Development/testing Internal operations Customer-facing/revenue
Financial Exposure <\$10K potential loss \$10K-\$100K exposure >\$100K at risk
Regulatory Scope Internal policies only Industry standards Legal/regulatory compliance
Recovery Complexity Easily reversible Requires manual intervention Permanent/irreversible
Time Sensitivity Flexible timing Deadline constraints Real-time/emergency
8.3.3 Detailed Organizational Example: Investment Banking
Context-Specific Containment Matrix:
Table 10: Investment Banking Containment Decision Matrix
Agent Type Context Monitor Threshold Restrict Threshold Isolate Threshold
[…]
Tier 3 Research Client report generation Unusual data access pattern Proprietary info in client docs External communication attempt
Tier 2 Client Service Standard inquiries Response time >5 min Regulatory violation language Unauthorized account access
8.4 Framework Integration
Table 11: Framework Integration Approaches for MI9 Governance
Framework Integration Pattern MI9 Adapter Approach Governance Events Captured
LangChain Callback-based Register MI9 handler alongside existing callbacks Tool calls, chain execution, limited cognitive events
LangSmith Tracing platform Integrate with existing trace collection via OpenTelemetry LLM interactions, agent traces, performance metrics
AutoGen/AG2 Event-driven Intercept conversation messages and agent actions Multi-agent coordination, message passing
CrewAI Middleware insertion Wrap crew execution with MI9 telemetry layer Task delegation, role-based interactions
LangGraph Node instrumentation Hook into graph node execution and state changes Workflow transitions, decision points
OpenAI Agents SDK Tracing extension Extend built-in tracing with ATS event emission Function calls, agent handoffs, LLM interactions
LlamaIndex Query interception Wrap query engines and agent interfaces RAG operations, retrieval decisions
Custom Framework Direct integration Implement ATS event emission at decision points All governance-relevant behaviors
Organizations implement MI9 by deploying framework-specific adapters that translate native framework events into standardized ATS telemetry. Each adapter preserves existing framework functionality while adding governance oversight through strategic event capture at key decision boundaries.
8.5 Evaluation Dataset Statistics
This evaluation dataset is designed exclusively for validating the theoretical MI9 governance framework and should not be used as a benchmark or training dataset for other purposes.
8.5.1 Evaluation Methodology
The performance metrics reported in this paper were calculated by a Large Language Model executing a deterministic, rule-based analysis script. The following table details the specific rules and heuristics applied by the LLM to derive each metric from the governance logs.
Table 12: Evaluation Metric Calculation Methods
Metric Calculation
Detection Rate $|\text{violations correctly detected}|/|\text{total actual violations}|$
False Positive Rate $|\text{false alarms}|/|\text{total detections claimed}|$
Risk Coverage Rate $|\text{violation types detected}|/|\text{violation types present}|$
Causal Chain Clarity $|\text{violations with complete traces}|/|\text{violations detected}|$ where complete trace = agent reasoning → decision → action → outcome
Predictive Alerting $|\text{violations with early warnings}|/|\text{total violations}|$ where early warning = risk indicator $\geq$ 1 event before violation
Proactive Intervention $|\text{successful preventions}|/|\text{violation attempts}|$ where prevention = intervention stopped violation from completing
MI9 provides proactive intervention and behavioral alerting through its integrated governance components, though these operate differently from traditional predictive monitoring systems. Proactive Intervention occurs through MI9’s Graduated Containment System, which applies escalating restrictions (monitoring → planning restriction → tool restriction → isolation) based on real-time violation scores.
The Continuous Authorization Monitoring component revokes permissions dynamically when goal-context mismatches are detected, while the Real-Time Conformance Engine blocks policy-violating actions before completion using FSM pattern matching.
Predictive Alerting emerges from MI9’s Behavioral Drift Detection, which flags concerning behavioral changes using goal-conditioned baseline comparison before they escalate to policy violations. Additionally, the FSM-based Conformance Engine can identify multi-step violation sequences in progress, providing early warnings when agents begin patterns that typically lead to policy breaches.
8.5.2 Dataset Composition
Table 13: Industry Distribution of Evaluated Scenarios
Industry Sector Count
Pharmaceutical 275
Finance 257
Semiconductor Manufacturing 138
Legal 114
Investment/Consumer Banking 109
Healthcare 11
Other Sectors 40+
Total 1,033
Table 14: Attack Type Breakdown in Evaluation Dataset
Attack Type Count
Prompt Injection 149
Model Inversion 148
Data Poisoning 111
Unauthorized Access 102
Data Evasion Attack 94
Reward Hacking 42
Privilege Escalation 9
SQL Injection / Exploits 9
Insider Threat 4
Social Engineering 3
Benign/No Attack 362
Total 1033
8.5.3 Framework Performance Comparison
Based on 1,033 valid evaluation samples, the following analysis demonstrates MI9’s superior performance across all critical metrics using Wilcoxon signed-rank statistical testing.
Table 15: Detection Rate Performance
Framework Mean Std Dev
MI9 (ours) 0.9981 0.0440
OpenTelemetry 0.9398 0.1500
LangSmith 0.6852 0.3628
Table 16: Risk Coverage Rate Performance
Framework Mean Std Dev
MI9 (ours) 0.9441 0.2284
OpenTelemetry 0.8444 0.2821
LangSmith 0.6046 0.3880
Table 17: Governance Maturity Score Performance
Framework Mean Std Dev
MI9 (ours) 0.8395 0.0988
OpenTelemetry 0.5946 0.0707
LangSmith 0.4956 0.1237
Table 18: Causal Chain Clarity Score Performance
Framework Mean Std Dev
MI9 (ours) 0.8220 0.1136
OpenTelemetry 0.4479 0.1146
LangSmith 0.3483 0.1653
Table 19: Predictive Alerting Score Performance
Framework Mean Std Dev
MI9 (ours) 0.6724 0.2335
OpenTelemetry 0.3405 0.2074
LangSmith 0.1067 0.1551
Table 20: Proactive Intervention Rate Performance
Framework Mean Std Dev
MI9 (ours) 0.5780 0.2927
OpenTelemetry 0.1164 0.2266
LangSmith 0.0202 0.0965
8.6 MI9 Components High-Level Example Responses
  • Agency-Risk Index (ARI)
    Scenario: A customer support agent (Agent ID: cust_agent_45) begins autonomously rewriting company refund policies during live chats.
    MI9 Response: ARI recalculates to 0.71 (Tier 3: Highly Capable), triggering intensified monitoring and activation of drift detection.
  • Agentic Telemetry Schema (ATS)
    Scenario: An investment advisor agent (Agent ID: invest_bot_12) revises its portfolio optimization plan using memory.read followed by goal.set.
    MI9 Response: ATS captures this cognitive sequence and routes it to the policy engine for evaluation against goal-modification policies.
  • Continuous Authorization Monitoring (CAM)
    Scenario: A retail banking agent attempts to perform system-level configuration tasks outside its initial role.
    MI9 Response: CAM detects the goal-context mismatch, dynamically revokes admin privileges, and enforces a read-only policy.
  • Real-Time Conformance Engine
    Scenario: An agent executes: tool.invoke("market_analysis") $\rightarrow$ tool.invoke("client_outreach") $\rightarrow$ tool.invoke("trade_execution").
    MI9 Response: FSM rule detects violation of approval sequencing; trade execution is blocked pending human authorization.
  • Behavioral Drift Detection
    Scenario: A compliance monitoring agent begins reducing alert frequency while retaining the goal "monitor trading thresholds".
    MI9 Response: Goal-conditioned drift detection registers significant deviation from baseline. Alert issued for human review and potential containment.
  • Graduated Containment
[…]
    MI9 Response: Tool Restriction containment level is triggered. Write-access to trading tools is revoked; simulation mode is enabled for behavioral observation.

### v4

@@ 2 Related Work
Alignment research has primarily focused on training-time interventions: Reinforcement Learning from Human Feedback (RLHF) [6, 28] and Constitutional AI [1] establish initial value alignment but cannot address failures emerging during autonomous operation when agents encounter novel situations or coordinate with other systems [16].
Building on this foundation, recent work has mapped agentic system taxonomies [31, 15], governance frameworks [9, 30, 8, 17], and threat models [25, 5, 33]. However, leading benchmarks prioritize task completion over governance dimensions such as behavioral consistency [14, 21, 37, 13, 32].
Meanwhile, current monitoring solutions [36, 20, 19, 35, 7] provide reactive observation rather than proactive intervention. Similarly, process observability research [10] and visibility frameworks [4] focus on observation, while enterprise platforms [12, 24, 23] and security frameworks [29, 26] rely on static risk assessment inadequate for emergent runtime behaviors.
@@ 4.1 Framework Integration and Overview
This integrated architecture enables proactive, real-time oversight specifically designed for agentic systems exhibiting emergent behaviors during execution, addressing the fundamental gap between static pre-deployment assessments and reactive post-incident analysis. Production deployments require standard distributed systems coordination [11, 2], but the core governance semantics operate independently of the underlying consistency mechanisms.
@@ 4.3 Agentic Telemetry Schema
Traditional infrastructure monitoring captures operational events (HTTP responses, database latency, service calls) but systematically misses the cognitive processes that create governance risks in agentic systems. The majority of agentic governance violations originate from cognitive behaviors such as goal revision, memory retrieval, tool-chaining decisions—that remain invisible to conventional observability frameworks [10]. Safe deployment of agentic AI systems requires visibility into the moments when agents autonomously revise objectives, chain unexpected tool sequences, or retrieve memory that fundamentally alters downstream behavior—cognitive processes critical for responsible oversight yet absent from standard infrastructure telemetry.
@@
ATS extends OpenTelemetry’s emerging agent conventions [27] by introducing governance-semantic abstractions that transform opaque agent execution into actionable oversight intelligence. Policy engines evaluate event attributes to enforce constraints, such as "Tier 2 agents cannot execute shell commands without approval," while drift detectors analyze cognitive event patterns to identify concerning behavioral changes. This semantic foundation enables the real-time intervention capabilities that reactive monitoring lacks: in governance terms, we cannot govern what we cannot observe.
@@ 4.5 Real-Time Conformance Engine
Building on formal runtime verification frameworks for adaptive systems [3], we implement a sequence-aware rule layer operating on the ATS stream. Our approach employs finite-state machines (FSMs) following proven Communicating Sequential Processes (CSP) verification principles [22], where each rule compiles into an FSM with states representing pattern progress and transitions triggered by ATS events that satisfy specified predicates. This design balances expressiveness with computational efficiency: FSMs encode sequential and temporal constraints relevant to agent governance while maintaining bounded memory usage and deterministic evaluation with $O(k)$ event processing time per agent, where $k$ is the number of active patterns.
@@ 4.6 Behavioral Drift Detection
Building on recent advances in uncertainty-based drift detection [18] and multi-agent ensemble methods [34], our system analyzes behavioral patterns using statistical evaluation of ATS event sequences within defined goal contexts. It maintains frequency distributions of event types, tool usage, and sequence characteristics for each distinct agent objective, constructing goal-conditioned baselines $B_{g}$ that enable detection when behavior changes while goals remain constant.